Ms. Masungsong is the sole owner of the Narnia Company. Presented below are the items to be consolidated for the immediate preparation of the Statement Of Financial Position, for the year 2020. 


Long-term portion of bonds payable 400% of Owner’s Withdrawals 

Property, plant, and equipment 30% of Total Assets

 Notes Receivable ₱ 500, 000 

Comprehensive Income 200% of Accounts Payable 

Raw materials Inventory 50% of PPE

 Prepaid Rent Expense 85, 000 

Cash and Cash equivalent 20% of Total Assets

 Owner’s Withdrawals 20% of Owner’s Capital

 Accounts Payable 60% of Owner’s Capital

 Unearned Revenue 480, 000 

Owner’s Capital 3, 000, 000 

Long-term Receivable 80% of raw materials inventory 

Long-term Investments ? 

Notes Payable 850, 000



a. Prepare the SFP using the ACCOUNT FORM 

 b. Prepare the SFP using the REPORT FORM 




1
Expert's answer
2021-09-12T19:32:32-0400

Solution:

Derive missing accounts figures:

Owners Capital = 3,000,000

Owners withdrawal = 20"\%" of owners capital = 20"\% \times" 3,000,000 = 600,000

Long-term portion of bonds payable = 400"\%" of Owners withdrawal = 400"\% \times" 600,000 = 2,400,000

Accounts Payable = 60"\%" of owners capital = 60"\% \times" 3,000,000 = 1,800,000

Comprehensive income = 200"\%" of Accounts payable = 200"\% \times"1,800,000 = 3,600,000

Total Liabilities & Owners Equity = 11,530,000

Since Total Liabilities & Owners Equity is normally equal to Total Assets.

Therefore, Total Assets = 11,530,000

Cash & Cash Equivalents = 20"\%" of Total Assets = 20"\% \times"11,530,000 = 2,306,000

Property, Plant and Equipment = 30"\%" of Total Assets = 30"\% \times"11,530,000 = 3,459,000

Raw Materials Inventory = 50"\%" of PPE = 50"\% \times"3,459,000 = 1,729,500

Long-term receivable = 80"\%" of raw materials inventory = 80"\% \times"1,729,500 = 1,383,600

 

The SFP calculations using the account form and report form are as follows: